OpenStudy (anonymous):

use double-angle identities to solve [0,2pi] -4sinx.cosx = 0.5

OpenStudy (anonymous):

i'm not so sure it is right or not

OpenStudy (anonymous):

but when I put it in the answer box, it tells me wrong :(

OpenStudy (tkhunny):

It's pretty subtle, but there is one tiny thing. You may have used the rounded values from the first two answers, and a rounded value of \(\pi\) to create the second two answers. \(\theta_{1} = \dfrac{\pi - asin(-0.25)}{2} = 1.697136454365935946\) Rounded 1.6971 \(\theta_{2} = \dfrac{2\pi + asin(-0.25)}{2} = 3.0152525260187539117\) Rounded 3.0153 \(\theta_{3} = \dfrac{\pi - asin(-0.25)}{2} + \pi = 4.8387291079557291844\) Rounded 4.8387 \(\theta_{4} = \dfrac{2\pi + asin(-0.25)}{2} + \pi = 6.1568451796085471502\) Rounded 6.1568 Round later and never use a short rounded number to create subsequent values.

OpenStudy (tkhunny):

Note: Amazingly clear and legible work!! Keep that up!